Gyne Company makes Product X using special plastic as the main material. The company made 1,000 units of X using 4,500 lbs of plastic. The plastic actually cost the company $22 per lbs (total $99,000). According to the company’s standards, each unit of X should require 5 lbs of plastic, at a price of $20 per lbs.

(a) What cost of plastic should have been incurred to make 1,000 units of X?

(b) How much is the spending (total) variance

(c) Is the spending variance favorable or unfavorable?

            (d) Compute the price variance

            (e) Is the price variance favorable or unfavorable?

            (f) Compute the quantity variance

            (g) Is the quantity variance favorable or unfavorable?
Answer #1

(a) Cost of plastic that have been incurred to make 1000 units of X:

Given, To produce 1000 units, company requires 4500lbs of plastic

Actual Cost of plastic = $22 per lbs

Therefore, total cost of plastic to make 1000 units of X
= 4500 lbs × $22
= $99000

(b) Calaculation of Spending varaince:

= Actual costs of materials - standard cost of materials
= $99000-(standard quantity of materials×standard rate)
= $99000 - (5000lbs × $20)
= $99000 - $100000
= $1000 U

*Standard qunatity of materials = 1000 × 5 lbs = 5000 lbs

(c) In case of soending variance if Actuals costs < standard costs, The spending Variance is Unfavourable.

(d) Calaculation of price variance

= (Actual price - standard price)× Actual quantity
= ($22 - $20) × 4500
= $2 × 4500
= $9000 U

(e) The price variance is Unfavourable, since Actuals > standards

(f) Calaculation of Quantity Varaince

= (Actual quantity - standard qunatity)×standard price
= (4500 - 5000) × $20
= -500 × $20
= $10000 F

*Standard quantity = 1000 × 5lbs = 5000 lbs

​​​​​​(g) The quantity variance is Favourable, since Actuals < standards

______×______

Do Vote, All the best,

Happy Chegging.